### v1

# Advancing Compositional Awareness in CLIP with Efficient Fine-Tuning

Abstract
Vision-language models like CLIP have demonstrated remarkable zero-shot capabilities in classification and retrieval. However, these models often struggle with compositional reasoning – the ability to understand the relationships between concepts. A recent benchmark, SugarCrepe++ [10], reveals that previous works on improving compositionality have mainly improved lexical sensitivity but neglected semantic understanding. In addition, downstream retrieval performance often deteriorates, although one would expect that improving compositionality should enhance retrieval. In this work, we introduce CLIC (Compositionally-aware Learning in CLIP), a fine-tuning method based on a novel training technique combining multiple images and their associated captions. CLIC improves compositionality across architectures as well as differently pre-trained CLIP models, both in terms of lexical and semantic understanding, and achieves consistent gains in retrieval performance. This even applies to the recent CLIPS [30], which achieves SOTA retrieval performance. Nevertheless, the short fine-tuning with CLIC leads to an improvement in retrieval and to the best compositional CLIP model on SugarCrepe++. All our models and code are available at https://clic-compositional-clip.github.io.
1 Introduction
Recent advances combining multiple modalities via large models have propelled us toward systems with greater capabilities. In the vision-language domain, models like CLIP [40], BLIP [22], FLAVA [22] and ALIGN [18] have been particularly transformative, by training on image-text paired datasets. These Vision-Language models (VLMs) showcase not only enhanced performance on native tasks like retrieval, captioning, etc., but also enable remarkable zero-shot capabilities in image classification [24, 30], segmentation [51], human-level perception of images [12], and similar other tasks. This suggests that CLIP like models are good at associating individual concepts with images. However, several works have shown [49, 46, 10] that CLIP models struggle in understanding how concepts combine to form complex meanings, i.e., compositionality.
As initially highlighted in [49], these models tend to learn a “bag of words” representation of the multimodal data, making them incapable of solving simple compositional tests. For instance, when given an image of “a man in a red shirt next to a gray horse", CLIP models have been shown to choose “a man in a gray shirt next to a red horse" as the more similar of the two captions (see [10] for more such examples). To this end, several new benchmarks (WinoGround [46], VALSE [35], Crepe [32], SugarCrepe [17], SugarCrepe++ [10]) have been created to test compositionality in VLMs.
While SugarCrepe [17] mainly tests whether a model can distinguish between lexically similar but semantically different texts (: “The dog is chasing a cat” vs. : “The cat is chasing a dog”), it has been pointed out in SugarCrepe++ [10] that this is not sufficient for true understanding. Given an image of a dog chasing a cat, models should attach higher similarity of the image to both and a semantically equivalent but lexically dissimilar text (“The cat is being chased by a dog”) than to the wrong text, . Surprisingly, it turns out that improving compositionality on SugarCrepe++ is non-trivial. As can be seen in Table 1, previous works perform very well on SugarCrepe (e.g., DAC [8]), but their performance on SugarCrepe++ is even worse than the pre-trained model.
Since capturing semantic similarity is crucial for the downstream performance of CLIP models, in particular retrieval, we focus on the evaluation of compositionality with SugarCrepe++.
In this paper, we introduce Compositionally-aware Learning in CLIP (CLIC). This technique fine-tunes only the text-encoder in CLIP models by leveraging already available high-quality captioned datasets like PixelProse [44] or common text-image datasets like Laion [41], which we recaption using CogVLM [47]. Using a novel technique of combining images and captions allows us to generate positives as well as hard-negatives with minimal additional overhead, e.g., we do not require an LLM for generating hard-negatives [8], nor do we need to generate synthetic images [37]. In Figure 1, we show that CLIC consistently improves both on SugarCrepe++ and on image-to-text and text-to-image retrieval benchmarks compared to the pre-trained CLIP-ViT-B/32 model [40], whereas existing methods do not achieve consistent gains. To summarize, our contributions are as follows.
- 1.
2.
We use CLIC to improve CLIPS [30], a recent version of CLIP trained to enhance retrieval performance. Our fine-tuned model, CLIPS + CLIC, yields SOTA numbers for CLIP-like models, with a average improvement on Image-to-Text (ITT) set of SugarCrepe++ and even improving on their already SOTA results in text- and image-retrieval, .
2 Methods and benchmarks for improving compositionality
Recent studies on VLMs have revealed significant limitations in their compositional abilities [49, 17, 46, 32] via specially designed benchmarks. Throughout these works, several types of image-text models have been shown to lack compositional reasoning, ranging from those using just contrastive loss (CLIP [40], LaCLIP [11], SigLiP [50]) to the ones using different losses or architectures like CoCa [48], FLAVA [42], CLIPA [29], BLIP [15]. In this section, we introduce our notation for standard contrastive learning and provide details on methods to improve compositionality and the associated benchmarks.
2.1 Contrastive learning and baselines
The contrastive loss is the backbone for training VLMs like CLIP. In contrastive training, image-text paired data is embedded into a joint embedding space via the text and vision encoders. Formally, let and define the normalized embeddings of the vision- and text-encoder of the CLIP model. Then, given a batch of image-caption pairs, the model is trained using an average of image-to-text and text-to-image contrastive losses,
Since training batches rarely include hard negative samples that challenge the “bag of words" representation, previous methods add hard negatives by slightly modifying captions to change their meaning, encouraging the model to recognize compositional cues. Denoting by the hard-negative augmentation of the caption ; methods such as NegCLIP [49] and TripletCLIP [37] adapt the image-to-text loss as follows,
2.2 Compositionality enhancing methods
Using hard negatives to enhance compositionality in VLMs was first introduced in NegCLIP [49], which employs spaCy [16] to swap words and phrases within a sentence. Similarly, SVLC [9] generates hard-negatives with two configurations: (i) SVLC-R, which uses a rule based generation with a lookup table and (ii) SVLC-R+L, which additionally uses an LLM based masked input completion, where masking is done for specific concepts using spaCy [16].
Dense and Aligned Captions (DAC) [8] uses the same method to create hard negatives, but first generates higher quality captions by passing the image through BLIP-2 [23]. In DAC-LLM, the enhanced captions are used as input to GPT-Neo-2.7B [13] for detailed and dense captions, and in DAC-SAM, segments of the original image, created using SAM [19], are captioned via BLIP-2. The model is then fine-tuned using a combination of (Eq. 1), (Eq. 3) and MIL-loss adapted from [34], which uses multiple negatives for each image.
TripletCLIP [37] uses rewritten captions from LaCLIP [11] and utilizes an LLM to generate hard-negatives by changing concepts such as objects, attributes, and relationships that change the meaning of the original caption while slightly distorting it. Afterward, a synthetic image is generated for the hard-negative captions using a text-to-image diffusion model. Finally, training is done by using a sum of image-text and text-image based from Eq. 2.
2.3 Compositionality benchmarks
Several benchmarks have been proposed to measure the lack of compositionality in VLMs. These benchmarks usually add a negative caption (semantically different) and measure cross-modal Image-to-Text (ITT) similarities across different captions. Among them is WinoGround [46], which uses two images and two corresponding captions that differ only in word order. The task of the model is to assign to each image the corresponding text and vice versa. However, it has been shown in [7] that models might fail on this benchmark for reasons other than compositionality, such as the ability to locate small and out-of-focus objects. SugarCrepe [17] measures image to text similarity and is built on the MS-COCO validation set. For each image, it requires the model to identify the correct caption from a pair of lexically similar, but semantically different captions ( and ).
Recently, SugarCrepe++ [10] extended this idea by introducing , an additional caption that is semantically similar to , but lexically different from both and . To answer correctly, the model needs to assign a higher score to and compared to the score assigned to . This makes SugarCrepe++ harder and more meaningful than SugarCrepe, as the model can no longer rely solely on lexical differences but has to instead understand the semantic differences between the captions. Therefore, we predominantly focus on SugarCrepe++ in this work.
On top of ITT, SugarCrepe++ also measures the uni-modal Text-to-Text (TOT) scores between the original, the semantically different, and the lexically different captions. The TOT is interesting for measuring the compositionality of the text encoder, but the ITT is arguably more important for the cross-modal downstream tasks for which CLIP models are used. Nonetheless, we report TOT for all experiments conducted in this work. Furthermore, we would also like to remark here, that many methods like [49, 43] train on MS-COCO, which may lead to knowledge leakage, as the captions and images in benchmarks like SugarCrepe and SugarCrepe++ come from the same data source.
3 Compositionally-aware learning in CLIP
Large-scale web-crawled datasets are known to be noisy, with many captions not reflecting their corresponding images [25, 3, 51]. This can be detrimental to compositional learning, which relies on minor modifications that alter semantic meaning. To improve the compositionality, we opt for captions that better align with the images and describe the images in detail. For this, we use approximately sample subsets taken from different data sources, like, RedCaps and CC12M from PixelProse [44], and our CogVLM [47] recaptioned Laion. Details about the datasets can be found in Appendix B. We train several versions of CLIC using each dataset to demonstrate the generalizability of our approach across data sources. Additionally, since some baselines are trained using MS-COCO, for a fair comparison, we train another version on this dataset, the details of which are in Section 4.
3.1 Generation of positives and hard-negatives: concatenation of images and captions
To overcome the “bag of words" limitation, we aim for the model to distinguish between captions that are lexically similar but semantically different. In order to achieve this, we construct hard negatives that retain the same words in the caption but no longer correctly describe the image. For efficiency and diversity, we create these examples by concatenating pairs of images and swapping words across their two captions, ensuring that the resulting caption does not represent the concatenated images. To help the model learn that different lexical or syntactic sentences can represent the same image, we create multiple positives. For this, we use dense captions that provide varied descriptions of each image. The concatenation enables the number of combinations to grow quadratically with the dataset size, allowing CLIC to generalize better than other methods, see discussion in Section B.3. We detail our data preparation strategy based on concatenating random image-caption pairs in the following section. The process is illustrated in Figure 2 and described in Algorithm 1 in Appendix C. See Section B.4 for a discussion on a non-random approach.
During training, in each iteration, we sample a batch of image-caption pairs, . We then randomly select additional pairs , and ensure that each selected image has the same orientation as . We denote by the process of concatenating two images or sentences and by the process of randomly shuffling the order of the two images or sentences before concatenation. We start by concatenating the images in a random order to mitigate potential biases and term the new image . Then, from the two associated captions and , we create five new captions for the image – four positive captions and one negative. As each original caption consists of several sentences, in the following, we denote by the sentence of the caption.
: We concatenate the first sentence of each caption and term this caption . This choice is based on the observation that the first sentence of a caption generally provides a high-level description of the image, whereas subsequent sentences describe more specific details, as demonstrated in the examples in Appendix B.1.
: We shuffle the order of the two sentences . We do this as both concatenated captions describe the corresponding concatenated image, irrespectively of the order of the sentences; thus, the model should be invariant to the order of the sentences.
: We then sample (without repetition) two additional sentences from each caption and concatenate them in a random order , and for random indices . This step reinforces the model’s ability to distinguish incorrect captions from correct ones, even when the incorrect caption presents a general description of the image while the correct caption describes only a specific part. This is different from previous works, such as DAC [8], where multiple positives exist, but each caption faces its own hard-negative. An ablation on the number of positives can be found in Table 13.
: Next, we leverage the spaCy package [16] to generate hard-negatives. We decompose the two first sentences into individual words and identify their respective linguistic categories using spaCy. A table detailing these categories, along with examples, is provided in Table 5 in Appendix B. To construct hard-negatives, we randomly select a linguistic category that is common to both sentences and choose two words from that category – one from each image. If no common category exists, we randomly select a word from each image and swap them. Certain categories, such as punctuation, are excluded from this process (see Appendix B.2 for details). To prevent over-fitting to specific benchmarks, we do not impose constraints on the selection of particular categories or word positions. As a result, the generated sentence no longer accurately represents the concatenated image unless, by chance, the swapped words share the same meaning.
3.2 Training of CLIC
The concatenated image is fed into the model along with the five captions. We employ a combination of different loss functions, each targeting a specific desired property of the model.
Contrastive loss: The standard CLIP loss (1) extended to handle the four positive captions,
where is the positive caption of an image for and .
Hard-negative loss: We compute a hard-negative loss between each of the positive captions and the negative caption of the same image, similar to Eq. 3,
where is the hard-negative text of . We use a separate loss for the hard-negatives to ensure they always influence the training and are not “masked” by other terms in the standard clip loss.
Uni Modal loss: The -distance between the first positive and its shuffled version :
Since the first two positives are simply shuffled versions of each other and the concatenated images are unrelated, there is no inherent difference between the positives (except for biases in the pre-trained model). Thus, we use this loss to teach the model to be invariant to syntactic alterations that do not change the meaning of the texts. This approach differs from the text-text similarity loss used in SVLC, where the loss is between two different descriptions of the same image, one of which was generated by a language model.
The final objective function we use is a weighted combination of these three losses:
where , , and are hyperparameters controlling the contribution of each loss term. To prevent the model from deviating too much from the pretrained CLIP when training only on concatenated images, we employ the standard clip training every second iteration, using a single image, the first sentence, and the standard clip loss as in Eq. 1.
SugarCrepe++ [10] | WinoGround [46] | SugarCrepe [17] | ||||||||
Replace | Swap | Text | Image | Group | Add | Replace | Swap | |||
Method | ITT | TOT | ITT | TOT | Score | Score | Score | ITT | ITT | ITT |
CLIP [40] | 69.5 | 60.5 | 45.7 | 25.9 | 31.2 | 11.0 | 8.7 | 72.9 | 80.0 | 62.7 |
NegCLIP [49] | 70.5 | 74.1 | 56.4 | 44.8 | 30.2 | 11.0 | 8.0 | 87.6 | 85.4 | 76.5 |
DAC-LLM [8] | 53.7 | 59.6 | 32.2 | 18.1 | 22.5 | 10.5 | 4.7 | 93.7 | 89.4 | 74.6 |
DAC-SAM [8] | 52.3 | 60.2 | 30.7 | 18.4 | 21.2 | 12.2 | 7.5 | 91.5 | 87.0 | 73.6 |
SVLC-R [9] | 64.0 | 69.5 | 51.4 | 28.4 | 25.2 | 8.5 | 6.7 | 85.3 | 78.9 | 68.8 |
SVLC-R+L [9] | 61.8 | 70.0 | 45.9 | 26.4 | 28.0 | 8.2 | 5.7 | 78.4 | 75.8 | 65.4 |
CoN-CLIP [43] | 69.0 | 70.8 | 48.0 | 32.5 | 29.5 | 10.0 | 7.2 | 83.0 | 79.7 | 65.3 |
TripletCLIP [37] | 73.5 | 72.7 | 43.4 | 33.2 | 30.7 | 11.2 | 8.2 | 86.6 | 88.0 | 69.6 |
CLIC-COCO | 73.5 | 79.4 | 52.9 | 47.2 | 31.0 | 13.5 | 10.0 | 85.5 | 83.8 | 69.3 |
CLIC-LAION | 75.6 | 60.1 | 61.1 | 27.9 | 31.5 | 11.5 | 9.2 | 84.5 | 84.0 | 73.7 |
CLIC-CC12M | 74.4 | 62.5 | 60.6 | 30.0 | 31.0 | 11.8 | 9.5 | 89.7 | 85.3 | 74.3 |
CLIC-RedCaps | 76.0 | 57.6 | 61.5 | 23.1 | 32.2 | 12.2 | 10.7 | 86.5 | 84.8 | 72.6 |
3.3 Differences between CLIC and other methods
Other approaches for generating hard text negatives differ from ours in computational cost and methodology. In DAC [8] and SVLC [9], hard-negatives tailored explicitly for specific compositionality tasks (e.g., replacing specific categories like colors, actions, etc.) are either found through look-up tables or created with computationally expensive LLM-based generation. Similarly, TripletCLIP uses an LLM to generate hard-negatives; for details, see Section 2.2. Closest to our work is NegCLIP, which also uses word swapping. However, NegCLIP focuses on specific word classes such as adjectives, adverbs, verbs, and nouns as described in Appendix D. We allow a swap of any two random words as long as they are from the same class, e.g., prepositions, numbers, and pronouns, as seen in Section B.2. Our approach offers two main advantages: cheaper computational cost compared to LLM-based generation and a design that doesn’t target any specific benchmark, reducing the risk of over-fitting on existing evaluation benchmarks. Other methods also add hard-negative images, further increasing overall computational complexity. NegCLIP and CoN-CLIP search for hard images to add to the batch, and TripletCLIP generates hard images using a diffusion model. In addition, none of the methods use a combination of image pairs and corresponding text descriptions as in our method.
Fine-Tuning | Classification | Text-Ret. | Image-Ret. | ||||
Method | Data | IMNET | ZS-10 | COCO | F30k | COCO | F30k |
CLIP [40] | - | 63.3 | 61.4 | 74.1 | 95.1 | 54.6 | 83.5 |
NegCLIP∗[49] | MS-COCO | 60.9 | 60.7 | 83.6 | 96.4 | 72.2 | 91.7 |
DAC-LLM [8] | CC3M | 51.1 | 54.3 | 63.3 | 79.3 | 58.1 | 88.3 |
DAC-SAM [8] | CC3M | 52.3 | 53.9 | 57.3 | 85.6 | 58.6 | 82.2 |
SVLC-R [9] | CC3M | 58.8 | 58.4 | 70.4 | 93.0 | 61.1 | 87.0 |
SVLC-R+L [9] | CC3M | 59.7 | 59.0 | 68.9 | 90.4 | 61.1 | 87.0 |
CoN-CLIP∗[43] | CC3M,MS-COCO | 63.2 | 61.3 | 71.4 | 90.7 | 55.5 | 84.5 |
TripletCLIP [37] | CC3M,CC12M | 54.8 | 52.3 | 72.3 | 91.5 | 56.8 | 85.1 |
CLIC-COCO∗ | MS-COCO | 62.7 | 60.7 | 82.9 | 97.1 | 68.2 | 90.2 |
CLIC-LAION | Laion | 61.7 | 61.0 | 75.9 | 95.0 | 60.0 | 86.7 |
CLIC-CC12M | CC12M | 62.2 | 60.6 | 76.9 | 95.8 | 60.8 | 87.9 |
CLIC-RedCaps | RedCaps | 62.3 | 60.2 | 76.0 | 95.6 | 59.5 | 86.3 |
4 Experiments and evaluation
Here, we describe the experimental details and the main results of our work.
4.1 Experimental setup
Dataset.
For CLIC, we fine-tune models either with CogVLM [47] recaptioned Laion images, or PixelProse [44] recaptioned images from RedCaps and CC12M (see Appendix B for details).
We stress that this approach keeps our training data independent of MS-COCO, which is the basis for benchmarks like SugarCrepe, SugarCrepe++, and retrieval evaluations. By that, we
ensure CLIC always functions as a zero-shot model.
However, to fairly compare with the ViT-B/32 model of NegCLIP, which is trained on MS-COCO, we also fine-tune one model on the MS-COCO [27] dataset.
In our experiments, we denote each model by the respective dataset it was trained on, i.e., CLIC-COCO, CLIC-LAION, CLIC-RedCaps, or CLIC-CC12M.
Models.
To be comparable with previous compositionality enhancing methods, we base the majority of our experiments and the ablations on ViT-B/32 models pre-trained by OpenAI [40].
However, to highlight the adaptability and scalability of our approach, we also fine-tune other commonly used pre-trained architectures, such as ViT-B/16 and ViT-L/14 from [40]. Up to our knowledge, among methods improving compositionality, only CoN-CLIP [43] provides checkpoints for these architectures.
Additionally, we fine-tune more recent ViT-L equivalent VLMs, including CLIPS and CLIPA to show that our method can enhance compositionality even in models that already possess compositional abilities and are pre-trained on varying datasets such as Re-DataComp (CLIPS), WIT (CLIP) and DataComp (CLIPA). More model specific details can be found in Appendix D.
Training details.
In all of our experiments for CLIC, we keep the vision encoder frozen and fine-tune only the text encoder. An ablation showing that this works better than training the entire model can be found
in Table 12 in the appendix.
We find it really intriguing that the main improvement in terms of compositionality can be achieved by just fine-tuning the text encoder.
We defer other training details like LR, loss parameters in Eq. 7, etc, and a discussion on our NegCLIP baselines (denoted everywhere with a †), used whenever there are no available checkpoints to Appendix D.
Evaluation.
We evaluated the models on the compositionality benchmarks SugarCrepe++ [10], WinoGround [46], and SugarCrepe [17]. Additionally, we assess the effects of fine-tuning on downstream tasks such as classification on ImageNet [5] and on ZS-10, which computes the average score across ten standard classification datasets, detailed in Section D.2. In addition, we report image () and text retrieval () Recall@5 scores on MS-COCO (Val2017) [27] and Flickr30k [38]. We show Recall@1 numbers for retrieval and detailed SugarCrepe and SugarCrepe++ numbers for all the models in Appendix E.
4.2 Comparing to other methods
Given that most publicly available checkpoints of compositionality methods use ViT-B/32 based CLIP, we first focus on this architecture.
From Table 1, we observe that most methods perform worse than the pre-trained CLIP on SugarCrepe++. Notably, only NegCLIP, CoN-CLIP, and the CLIC models improve or are on par with the pre-trained baseline. Among these, CLIC-COCO is better than NegCLIP and CoN-CLIP (all three were trained on MS-COCO), while CLIC-RedCaps has lower TOT with a higher ITT (10.2% gain on pre-trained model across replace and swap). We also evaluate all models on the WinoGround benchmark, where, apart from CLIC, no other method improves on the pre-trained model, highlighting the effectiveness of CLIC. Moreover, CLIC improves in compositionality regardless of the fine-tuning dataset, showing that CLIC works for different sources of captions and images used for fine-tuning.
An interesting finding in Table 1 is that many of the models presented in the table perform well on ITT in SugarCrepe, but struggle for ITT in SugarCrepe++.
We remind the reader that ITT in SugarCrepe++ is an extended version of ITT from SugarCrepe. It additionally
includes lexically different but semantically similar positives, which the model should also favour over the negatives.
This stark difference in performance for methods like DAC (has SugarCrepe replace ITT of 89.4% but SugarCrepe++ ITT of only 53.7%) and SVLC suggests these models might not be truly learning compositional relations. More specifically, this suggests that although the models manage to detect lexical changes from the original text, they struggle when facing a lexically different text with the same meaning. The probable cause of this could be
the over reliance of these methods on using
very specific hard-negatives tuned for
previous benchmarks, see discussion in Section D.4.
Downstream tasks. We evaluate the models on standard retrieval benchmarks (MS-COCO and Flickr30k) and on classification tasks (ImageNet and ZS-10), as detailed in Table 2. By looking at both these tables, one sees that TripletCLIP performs well on SugarCrepe++ and WinoGround, but is found lacking in downstream tasks performance on ImageNet, ZS-10. We see that our CLIC models exhibit lower performance degradation in classification tasks compared to most models. In addition, CLIC-CC12M achieves the highest text retrieval performance among the models not trained on MS-COCO while also improving on the pre-trained model for image retrieval.
Downstream Evaluations | Sugarcrepe++ | Sugarcrepe | |||||||||
Method | Classification | COCO Ret. | Rep. | Swap | Add | Rep. | Swap | ||||
IMNET | ZS-10 | ITT | TOT | ITT | TOT | ITT | ITT | ITT | |||
CLIP [40] | 75.5 | 72.3 | 79.0 | 60.0 | 70.7 | 58.7 | 44.7 | 22.7 | 74.9 | 79.5 | 61.4 |
NegCLIP† | 73.9 | 70.6 | 77.7 | 65.9 | 70.6 | 63.5 | 50.5 | 35.8 | 83.6 | 82.8 | 70.0 |
CoN-CLIP∗[43] | 75.9 | 72.4 | 78.4 | 63.1 | 72.2 | 71.3 | 46.0 | 30.5 | 83.9 | 81.3 | 64.2 |
CLIC-LAION | 74.2 | 71.8 | 80.1 | 65.4 | 79.0 | 59.1 | 59.8 | 26.3 | 87.7 | 85.6 | 71.3 |
CLIC-RedCaps | 75.2 | 72.1 | 81.3 | 63.9 | 76.0 | 55.4 | 55.2 | 20.0 | 84.3 | 83.8 | 66.1 |
CLIP-A [24] | 79.6 | 78.3 | 85.4 | 70.5 | 74.8 | 74.8 | 45.8 | 30.8 | 85.8 | 82.9 | 63.8 |
CLIC-LAION | 78.4 | 78.0 | 84.8 | 71.5 | 78.4 | 66.3 | 55.1 | 30.7 | 92.9 | 88.2 | 71.6 |
CLIC-RedCaps | 78.9 | 78.3 | 85.2 | 71.2 | 80.2 | 62.8 | 61.2 | 22.8 | 93.4 | 85.9 | 71.5 |
CLIP-S [30] | 78.4 | 76.5 | 90.4 | 77.3 | 79.3 | 74.2 | 60.6 | 51.1 | 89.4 | 88.0 | 79.6 |
CLIC-LAION | 76.7 | 76.2 | 90.4 | 79.4 | 84.3 | 71.9 | 73.0 | 49.4 | 96.7 | 92.4 | 85.7 |
CLIC-RedCaps | 77.2 | 76.6 | 91.7 | 79.5 | 84.9 | 71.4 | 75.1 | 50.2 | 95.3 | 91.1 | 84.5 |
4.3 Ablation Study for CLIC
In Table 10 in Appendix E we ablate the components of CLIC in steps (C1-C5) and present the analogous non-concatenated, single image version of CLIC. The CLIC baseline (C1) is fine-tuning on the concatenated images with caption and in (C2) we add our single hard negative. These two steps are sufficient to improve on SugarCrepe but not on SugarCrepe++. This suggests that the hard-negative distractor helps the model to detect lexical changes, but not to distinguish between semantically similar sentences and ones with different meaning. Adding multiple positives (C3) and the uni-modal loss (C4) improves significantly on SugarCrepe++ while performance on SugarCrepe slightly reduces. Thus adding diverse positives and enforcing invariance of semantically identical captions lets the model identify semantically similar captions while being sensitive to lexically similar but semantically dissimilar captions. In the last step (C5) we add the iterates with single images and standard contrastive loss which further improves compositionality as well as downstream task performance by eliminating the potential bias of the model due to the concatenated images. The non-concatenated, single image variant of CLIC shows similar but smaller improvements in compositionality while being significantly worse in retrieval.
4.4 Generalization to larger architectures and models
Next, we measure the generalization of the method to larger architectures such as ViT-L/14 across models pre-trained on different datasets (CLIPA and CLIPS). Results on ViT-B/16 can be found in Table 16 in the appendix. We fine-tune each of the pre-trained models, CLIP, CLIPA, and CLIPS, on our 1M Laion subset and on RedCaps (PixelProse) subset using CLIC. As most other compositionality enhancing methods do not explore larger architectures, the only baselines we have are CoN-CLIP and our NegCLIP† (the details of which can be found in Appendix D). In Table 3, we can see that for all types of pre-trained models, CLIC-RedCaps shows consistent improvements on the ITT score for SugarCrepe++: +7.4% for CLIP, +9.4% for CLIPA, and +9.2% for CLIPS when averaged across replace and swap. These improvements are on a similar scale as the ones for the ViT-B/32 model, suggesting that even for larger architectures, CLIC can attain substantial gains.
SugarCrepe++ | WinoGround | |||||||||
Model | Params (M) | Replace | Swap | Average | Score | |||||
ITT | TOT | ITT | TOT | ITT | TOT | Text | Image | Group | ||
FLAVA [42] | 358 | 74.4 | 75.0 | 56.8 | 51.5 | 67.3 | 65.5 | 32.2 | 20.5 | 14.2 |
BLIP-Base [22] | 225 | 73.8 | 73.9 | 54.0 | 42.1 | 66.1 | 61.1 | 35.8 | 15.8 | 13.3 |
BLIP-Large [22] | 470 | 81.7 | - | 59.8 | - | 72.9 | – | 37.7 | 13.5 | 10.5 |
ViT-H/14 [41] | 986 | 73.8 | 72.1 | 48.8 | 39.1 | 63.7 | 58.8 | 33.5 | 12.7 | 10.5 |
ViT-bigG/14 [41] | 2540 | 76.6 | 73.2 | 51.5 | 40.2 | 66.6 | 59.7 | 35.5 | 15.0 | 12.0 |
CLIPS [30] | 427 | 79.3 | 74.2 | 60.6 | 51.1 | 71.9 | 64.6 | 36.2 | 16.0 | 12.5 |
CLIPS + CLIC | 427 | 84.9 | 71.4 | 75.1 | 50.2 | 81.0 | 62.9 | 41.7 | 17.5 | 15.2 |
What we find even more fascinating are the improvements for image and text retrieval attained by CLIC. Specifically, for CLIPS (which has to the best of our knowledge SOTA retrieval numbers of a CLIP model), CLIC further improves text retrieval by 1.3% and image retrieval by 2.2% on MS-COCO. These gains comes at a tiny fraction (0.01%) of the pre-training cost, as we only see 850k (RedCaps) resp. 1M (Laion) samples compared to 13B samples seen during pre-training of CLIPS.
4.5 Contrasting against SOTA
In Table 4, we pitch our best model in CLIPS + CLIC against the SOTA CLIP like and other models on SugarCrepe++. To the best of our knowledge, CLIPS + CLIC attains the best available ITT numbers on SugarCrepe++ across all architectures. We note that CLIPS in itself is already good for ITT on SugarCrepe++ but our short fine-tuning enhances it significantly further in terms of compositional understanding. For TOT, among CLIP like models CLIPS + CLIC performs similarly to the best model (CLIPS) in this category. We were unable to evaluate BLIP-Large for TOT, since the individual BLIP-Large encoders are not publicly available. The WinoGround results for CLIPS + CLIC are the best in text and group score, while the image score is the best among CLIP-like models. As far as we know, CLIPS + CLIC attains SOTA WinoGround scores among CLIP based models.
5 Conclusion
We introduce CLIC, a simple, low-cost, and scalable method for boosting compositionality in vision-language models, as measured by SugarCrepe++. This improvement comes with a minimal degradation in zero-shot classification, but enhanced text and image retrieval performance. CLIC works across different good caption quality datasets. We further illustrate the generalization of CLIC to variedly (dataset/training-scheme) pre-trained models and architectures. Lastly, we show that CLIC even improves performance on models already exhibiting strong compositional and retrieval abilities like CLIPS and yields SOTA results on SugarCrepe++ as well as image retrieval.
Acknowledgements
The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting AP and NDS. We acknowledge support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy (EXC number 2064/1, project number 390727645). We are also thankful for the support of Open Philanthropy and the Center for AI Safety Compute Cluster. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the sponsors.
Appendix A …Limitations
2.
Appendix B …Dataset preparation
3.
Appendix C …Algorithmic details
4.
Appendix D …Experimental details and discussions
5.
Appendix E …Additional experiments
Appendix A Limitations
Our work considers only a fraction of contrastive models, we believe CLIC should work similarly for other models like SigLIP [50], which we were unable to test due to compute and time constraints. Since all gains attained by CLIC are due to fine-tuning the text encoder, we cannot test CLIC fine-tuned models in large LVLMs like LlaVA [28] or VILA [26], which only use the vision encoder from CLIP.
Appendix B Data preparation
In this section, we describe the data preparation process in detail. Appendix B.1 showcases randomly selected images and captions from the PixelProse [44] dataset and from the Laion [41] dataset, along with captions generated for these images by CogVLM [47]. This part also presents further details on these datasets. In Section B.2, we explain the hard-negative generation procedure. In Section B.3 we show examples of the limitations of using a single image for generating negatives as done in [49]. Section B.4 shows ablations regarding other ways of combining images for CLIC.
B.1 Training datasets
In this section, we present our data generation process. To show that CLIC performs well across diverse datasets, we train models from three different sources. We took a subset of images from Laion with captions that were generated by CogVLM [47], using the query presented in Figure 3. We also used a recently proposed high-quality caption dataset in PixelProse [44], which contains available samples from the RedCaps subset. In order to compare faithfully to our baselines in Table 1 who use CC3M, we also use a 1.2M subset of CC12M that comes from the same data stream as CC3M to train a ViT-B/32 model. Since our method requires detailed captions, we use the recaptioned version of CC12M from PixelProse. These two datasets come from a different stream of internet data than Laion. A selection of random image-caption pairs from both kinds of datasets is provided in Figure 4. As can be seen from the examples, the first sentence in each caption usually provides a general description of the image, while the other part of the caption often captures more specific details. To clean the data, we removed the starting strings like “This picture depicts/shows/demonstrates:”. Then, since the generated captions tend to be longer than those in the standard datasets like Laion and DataComp (the average caption length in our recaptioned dataset is 78 words), we split the captions into individual sentences as done in [30, 51]. This results in an average sentence length of 16 words for captions generated by CogVLM [47]. We retain only images whose captions contain more than one sentence, which is more than of the 1M set for Laion. This approach favors captions that provide both general context and specific details about the images. For a fair comparison with NegCLIP, we also train CLIC with MS-COCO, the details of which are in Section 4.
Laion re-captioned images
An elderly man wearing glasses and a suit standing behind a lectern with a golden eagle design. He appears to be in a formal setting possibly delivering a speech or presentation. The background is dimly lit and there are some decorative elements such as flowers visible behind him.
A dimly lit cave or underground setting. The cave has various elements such as stalactites rocks and a chalkboard with some drawings on it. The atmosphere seems mysterious and adventurous suggesting that the characters might be on a quest or exploration.
A bar chart that represents the average sales price of homes in Centerville Georgia over a 12-month period from January 2010 to December 2010. The chart shows two sets of data: the average sales price of homes and the average sales price to list price ratio.
PixelProse images
In the foreground, there is a long bridge with a train on it. The bridge is made of metal and has a brown color. The train is red and white. Behind the bridge, there is a large body of water. The water is a blue-green color. There are some clouds in the sky. The image displays the Pamban Bridge, which is a still active train bridge located in Rameshwaram, India.
A pink and white frangipani flower. The flower is made up of five petals, with the pink petals on the outside and the white petals on the inside. The flower is surrounded by green leaves.
A small Bengal cat with brown fur and black stripes. The cat is lying on a cat tree, with its front paws hanging off the edge of the platform. The cat is looking at the camera with a curious expression.
B.2 Hard negative creation
POS Tag | Meaning | Example Words |
ADJ | Adjective | happy, blue, large |
ADP | Adposition (preposition) | in, on, under, with |
ADV | Adverb | quickly, very, well |
AUX | Auxiliary verb | is, was, will, do |
CCONJ | Coordinating conjunction | and, but, or |
DET | Determiner | the, a, an, this |
INTJ | Interjection | wow, oh, hey |
NOUN | Noun (common & proper) | dog, city, John |
NUM | Numeral | one, two, 100 |
PART | Particle | not, ’s (as in John’s) |
PRON | Pronoun | he, she, it, they |
PROPN | Proper noun | London, Microsoft |
PUNCT | Punctuation | ., !, ? |
SCONJ | Subordinating conjunction | because, although, if |
SYM | Symbol | $, %, & |
VERB | Verb | run, eat, write |
X | Other (foreign words, typos) | jdhfk, asdfg |
Our hard negative creation process utilizes the spaCy package to identify the part-of-speech (POS) of words in a sentence. We consider only the coarse-grained POS level, which is universal and not language-specific. The complete list of universal POS tags is provided in Table 5. For each sentence, we classify all words, group them by their POS, and determine eligible categories for swapping. A valid swap nomination requires at least two words from the same POS category, excluding the following: “AUX", “CCONJ", “DET", “INTJ", “PART", “PUNCT", “SCONJ", “SYM" and “X". Once a valid category is identified, we randomly select it and swap two randomly chosen words within that category. We deliberately did not select specific tags to avoid overfitting to current compositionally benchmarks. If no category falls into these cases, we swap two words at random.
B.3 In-scene vs across-scene swaps
Creating negatives by swapping words between sentences from different scenes offers several advantages over swapping within a single scene:
1.
Swapping words within a single sentence can lead to negatives due to the changes in word order. For example, the sentence from NegCLIP: “The horse is eating the grass" can become “The grass is eating the horse". However, swapping words across sentences from different images greatly increases the diversity of negative examples. Taking the example from the paper (Section 1), concatenating the image with an image of a dog chasing a cat may lead to the new negative: "The horse is chasing the cat. The dog is eating the grass". Since each image can be paired with each of the other images, this scales the diversity of negatives quadratically and helps the generalization of the method.
2.
Creating negatives from a single scene, like in NegCLIP, may result in a modified sentence that still accurately describes the image. This might happen as in many sentences, changing the order of the words does not change the semantic meaning of the sentence. This is different from changing words across scenes, where the concatenated sentence will reflect the concatenated image only if we by chance swap words that have the same meaning (note that we do not allow swaps of the same word, i.e., ‘man’ can not be swapped with ‘man’). Here are a few examples from the CC12M subset from PixelProse:
“In the top left corner, a person wearing a green jacket” – switching left and top does not change the meaning.
“there is a tuna salad with celery, onion, and mayonnaise” – changing onion and celery does not change the meaning.
“Bugs Bunny, Taz, Lola Bunny, and Daffy Duck are standing on either side” – changing Lola and Bugs.
“The Jeep has a black bumper, black wheels” – changing bumper and wheels.
“A dining room and living room" – changing dining and living.
“A black and gray backpack” – changing gray and black.
B.4 Non-random image concatenation
Combining images at random offers us several advantages such as computational efficiency and reducing the overfit to specific benchmarks. However, to test whether using related scenes might produce harder negatives and, by that, help us achieve better results (at the expense of higher computational resources) we combine images that share a common noun. To this end, we first extract all the nouns from the first sentence of each image using spaCy. Then for each image, we sample up to five images (without repetition) that share a common noun and use this to concatenate the images. We then continue as usual by swapping random words (excluding the common noun). Although we witnessed a higher loss (Eq. 5), indicating that the hard negatives are indeed harder, the results on the downstream tasks stayed roughly the same, as can be seen in Table 6
Downstream Evaluations | SugarCrepe++ | SugarCrepe | |||||||||||
Classification | Text Ret. | Image Ret. | Replace | Swap | Add | Rep. | Swap | ||||||
Method | IMNET | ZS-10 | COCO | F30k | COCO | F30k | ITT | TOT | ITT | TOT | ITT | ITT | ITT |
CLIP [40] | 63.3 | 61.4 | 74.1 | 95.1 | 54.6 | 83.5 | 69.5 | 60.5 | 45.7 | 25.9 | 72.9 | 80.0 | 62.7 |
CLIC-LAION | 61.7 | 61.0 | 75.9 | 95.0 | 60.0 | 86.7 | 75.6 | 60.1 | 61.1 | 27.9 | 84.5 | 84.0 | 73.7 |
Non-random | 61.7 | 60.7 | 75.9 | 95.0 | 60.0 | 86.6 | 75.0 | 60.1 | 61.8 | 28.5 | 84.3 | 83.3 | 73.0 |
Appendix C Algorithmic details
In this section, we present a pseudo-code of our algorithm, as discussed in Section 3 in the main paper. The pseudo-code can be found in Algorithm 1.
Appendix D Experimental details and discussions
In this section, we give a detailed description of various training and baseline design choices.
D.1 Further training details and discussions
As stated in the main paper, we train with alternating steps of CLIP loss (Eq. 1) and the proposed loss (Eq. 7). This decision is made to help the fine-tuned model retain its performance on downstream tasks, evident from the ablations in Table 10.
Computational resources.
All the work was carried out on A100 40G GPUs. The training runs per hour are across 4 GPUs. Running CLIC on smaller architectures like ViT-B/32 and ViT-B/16 took less than 30 minutes per GPU hour for the Laion dataset. Datasets such as RedCaps took longer due to high-resolution images. The larger ViT-L/14 models took around 1 GPU hour per dataset. The evaluation time is negligible, in the order of 20 minutes per model for all datasets on a single GPU.
Training parameters and details.
In the experiments, the loss parameters in Eq. 7 are set to , and . We train for one epoch on our 1M Laion subset and on our PixelProse dataset, while for MS-COCO, we trained for five epochs; ablation on the number of epochs can be found in Table 11. All experiments are conducted at an image resolution of . Our effective batch size is , and we use a cosine scheduler, where the warm-up phase is of the training time. The learning rate (LR) starts at , peaks at , and arrives at at the end. We used the standard AdamW [31] optimizer with beta parameters and set to with a weight decay of 0.1.
Training with COCO.
Although for most of the experiments, we use our 1M Laion subset or the PixelProse dataset, we also train with MS-COCO to be comparable to the original NegCLIP checkpoint from [49]. Note that due to training on MS-COCO all retrieval numbers on MS-COCO for NegCLIP from [49], and CLIC-COCO are no longer zero-shot. The same is also true to CoN-CLIP, who use COCO images as distractor in their method.
Our NegCLIP baselines.
To have a baseline for all models considered in this work, we train our own version of NegCLIP (denoted everywhere with †). It is used only when there are no available checkpoints (Table 3, Table 16 and Table 17), and in our ablation study (Table 12). This version of NegCLIP differs from the NegCLIP [49] as follows:
1.
We do not use hard negative images that are added to the batch.
2.
We use the same categories as in our method, while the NegClip paper employs six tags: two from the coarse-grained POS level (“ADV" and “ADJ"), two from the fine-grained level (“NN" for singular nouns and “NNS" for plural nouns), and two phrase-level categories (verb phrases and noun phrases).
3.
The vision encoder is frozen during training.
We use their negative loss calculation (see Eq. 2), where the hard negatives are part of the standard clip loss.
D.2 Zero-shot datasets
The 10 zero-shot classification datasets we use are a subset from the CLIP_benchmark111https://github.com/LAION-AI/CLIP_benchmark. Specifically, we use 1k images of each of the following datasets: Country-211 [39], Caltech-101 [14], OxfordPets [36], DTD [4], FGCV Aircrafts [33], StanfordCars [20], Cifar-10,100 [21], Food-101 [2].
D.3 Dataset from other methods
To improve compositionality, many methods use re-labelled/captioned versions of standard datasets. The details of these recaptions can be found in Table 7. TripletCLIP uses recpationed data from LaCLIP and DAC uses BLIP2 [23] to create better captions. Methods like CoN-CLIP and SVLC-R+L use the original captions and PaLM-2 and BERT [6] to generate hard-negatives for their base datasets.
Base | Re-caption | ||
Method | Source | Dataset | Method |
NegCLIP | [49] | MS-COCO | - |
DAC-LLM | [8] | CC3M | BLIP2 [23] |
DAC-SAM | [8] | CC3M | BLIP2 [23] |
SVLC-R | [9] | CC3M | - |
SVLC-R+L | [9] | CC3M | BERT [6] |
CoN-CLIP | [43] | CC3M | PaLM-2 [1] |
TripletCLIP | [37] | CC3M, CC12M | LaCLIP [11] |
CLIC-COCO | Ours | MS-COCO | - |
CLIC-LAION | Ours | Laion | CogVLM [47] |
CLIC-CC12M | Ours | CC12M | PixelProse [44] |
CLIC-RedCaps | Ours | RedCaps | PixelProse [44] |
D.4 Other methods do not generalize
From the results associated with SugarCrepe++ and SugarCrepe, one sees that other compositionality enhancing methods like SVLC-R, DAC etc., perform relatively well on SugarCrepe but do not generalize to SugarCrepe++. We believe this is due to the way these methods operate. Specifically, most of these methods create hard-negatives aimed at certain attributes like color, position, etc, which is exactly what benchmarks like ARO, and SugarCrepe (ARO addresses the same task, but SugarCrepe generates more fluent and sensical negatives) test for. For SugarCrepe++, in addition to the evaluation in SugarCrepe, an additional lexically different query text is introduced and since these methods do not account for general compositional/structural variations in their hard-negative, they fall short on SugarCrepe++. On the other hand, the random word swapping in CLIC makes our method focus not on the specific attributes but the general composition of different words in the query text. This makes our method generalize better and work equally well for both SugarCrepe++ and SugarCrepe.
Appendix E Additional experiments
In this section, we present more detailed results of the experiments in Section 4 in the main paper and provide additional experiments.
Extension of results from the main paper
As we only show averaged ITT and TOT numbers across different sets of SugarCrepe++ and SugarCrepe in the main paper, the full versions for ViT-B/32 models can be found in Table 8 and Table 9 resp. Similarly, for ViT-B/16 (omitted in the main part) and ViT-L/14, detailed results are shown in Table 16 and Table 17, respectively. Detailed compositionality results for the newer CLIP versions in CLIPA, EVA02-CLIP and CLIPS with their respective better versions achieved via CLIC can be found in Table 18. In Table 15, we give dataset wise accuracy of all the large models for the ZS-10 setting. WinoGround results for large model can be found in Table 19. The pre-trained checkpoints used for fine-tuning with CLIC are listed in Table 14.
Improvements across architectures
In Table 3 in the main part of the paper, we show how CLIC improves compositionality and retrieval for differently pre-trained models. We visualize these improvements in Figure 5. Fine-tuning with CLIC yields improvements on all models for ITT on SugarCrepe++ and SugarCrepe. The improvements in case of CLIP are as high as 7% and as much as 9% for CLIPS.
Ablating CLIC
In Table 10, we analyse the components of CLIC (), where we concatenate two images, along with its analogous version using only a single image (). For the single image version of our CLIC, fine-tuning on the Laion data () improves SugarCrepe results a bit and marginally improves for some cases of SugarCrepe++. Adding a single hard-negative (), drops the downstream performance without any improvement in compositionality. Adding more positives ( and ) improves ITT for SugarCrepe++ but leads to a decay in downstream and SugarCrepe numbers. Note that is identical to in this case, as one only has a single image. Adding the standard CLIP loss in recovers downstream performance back.
The ablations for CLIC reveals that starting from (the standard pre-trained CLIP model) and training with concatenated images and a single caption manages to improve some categories TOT in SugarCrepe++, but does not improve ITT scores on SugarCrepe++. Further adding a single hard-negative () does not consistently improve results on SugarCrepe++ but improves SugarCrepe results, suggesting that the hard-negative distractor might help the model to detect changes, but does not help the model to distinguish between semantically similar sentences to ones with different meanings. Furthermore, integrating the Uni-Modal loss function (6) () either improves or maintains performance in both ITT and TOT tasks, as it helps the model discover syntactic changes that do not change the meaning of the caption. In addition, since the model has been trained only on concatenated images, but standard benchmarks measure single-image tasks, we can see the recovery of the model during the standard CLIP loss iterations (, this is the CLIC-LAION version). Finally, our method demonstrates advantages over the single image-text () baseline (uses the additional positives and and the clip loss iterates).
Ablating training steps
In Table 11, we show how the proposed CLIC changes with more training steps. Here, we use MS-COCO for training which means 5 epochs corresponds to 2.5 Laion epochs of our dataset. From the table, we see increasing gains in SugarCrepe++ numbers when increasing the number of training step. However, this comes with a marginal degradation in zero-shot classification, which goes down from 63.2% to 62.7% on ImageNet as we move from 2 to 5 epochs. The same trend holds for our 1M set of Laion, where training for more than 1 epochs led to decay on ImageNet and ZS-10 numbers with marginal gains on SugarCrepe++. Hence, for our runs in the main part of the paper with Laion (CLIC-LAION) we only train for 1 epoch.
Ablating freezing of model components
In Table 12, we show how the proposed CLIC (frozen vision encoder) yields, on average, the best compositionality without sacrificing downstream performance. From the table, it is clear that freezing just the text encoder leads to the least amount of improvement on compositional benchmarks. Fine-tuning the whole model works better (in terms of retrieval) but the gains on SugarCrepe++ are smaller than freezing the vision encoder and it also degrades zero-shot classification performance. From these results, we infer that the most improving configuration is freezing the vision encoder, which is the final setting for CLIC. This also highlights that substantial gains can be made by improving the text encoder alone.
Ablating number of additional positives
In Table 13, we show the impact of adding additional positives for each concatenated image ( and from Section 3.1). Notably, increasing the number of positives has opposing effects on the TOT and ITT scores. In addition, after two positives, the effects on ITT in SugarCrepe++ become minimal, so we chose to add only two positives.
More retrieval results
In the main part of the paper, all results on MS-COCO and Flickr30k retrieval are reported with Recall@5. In Table 20, we additionally report Recall@1 for all ViT-B/32 models from Table 1. The overall ordering of methods is similar to Recall@5. Similarly, in Table 19, we additionally report Recall@1 for all ViT-L/14 models from Table 3.
Method | Replace-obj | Replace-att | Replace-rel | Swap-obj | Swap-att | AVG | ||||||
ITT | TOT | ITT | TOT | ITT | TOT | ITT | TOT | ITT | TOT | ITT | TOT | |
CLIP | 86.8 | 83.7 | 65.6 | 59.3 | 56.2 | 38.6 | 46.1 | 19.2 | 45.2 | 32.7 | 60.0 | 46.7 |
NegCLIP | 89.6 | 94.5 | 69.5 | 76.4 | 52.3 | 51.4 | 54.7 | 33.1 | 58.8 | 56.7 | 64.8 | 62.5 |
DAC-LLM | 65.7 | 76.8 | 47.7 | 59.5 | 47.6 | 42.3 | 31.4 | 11.4 | 32.9 | 24.8 | 45.1 | 43.0 |
DAC-SAM | 64.3 | 75.8 | 43.9 | 56.1 | 48.6 | 48.6 | 27.7 | 11.4 | 33.6 | 25.4 | 43.6 | 43.5 |
SVLC-R | 82.9 | 89.5 | 61.5 | 67.4 | 47.6 | 51.5 | 49.4 | 20.4 | 53.3 | 36.3 | 58.9 | 53.0 |
SVLC-R+L | 80.9 | 91.6 | 57.1 | 67.0 | 47.3 | 51.3 | 42.8 | 18.4 | 48.9 | 34.4 | 55.4 | 52.5 |
CoN-CLIP | 88.1 | 91.5 | 66.6 | 69.4 | 52.3 | 51.6 | 44.1 | 22.4 | 51.9 | 42.6 | 60.6 | 55.5 |
TripletCLIP | 86.8 | 92.3 | 71.7 | 74.0 | 61.9 | 51.7 | 38.8 | 24.1 | 47.9 | 42.3 | 61.5 | 56.9 |
CLIC-COCO | 90.2 | 97.3 | 72.3 | 81.6 | 58.0 | 59.3 | 48.6 | 33.5 | 57.2 | 60.8 | 65.3 | 66.5 |
CLIC-LAION | 90.0 | 84.5 | 75.8 | 57.9 | 61.0 | 38.0 | 61.2 | 22.4 | 61.0 | 33.4 | 69.7 | 47.2 |
CLIC-RedCaps | 90.1 | 84.3 | 75.1 | 52.4 | 62.9 | 36.3 | 60.8 | 17.9 | 62.4 | 28.2 | 70.3 | 34.6 |
Methods | ImageNet | Add | Replace | Swap | ||||
Obj. | Att. | Obj. | Att. | Rel. | Obj. | Att. | ||
CLIP | 63.3 | 77.2 | 68.6 | 90.9 | 80.1 | 69.1 | 61.2 | 64.3 |
NegCLIP | 60.9 | 88.8 | 84.7 | 93.8 | 87.7 | 73.9 | 75.5 | 76.4 |
DAC-LLM | 51.1 | 89.6 | 97.7 | 94.4 | 89.3 | 84.4 | 75.1 | 74.2 |
DAC-SAM | 52.3 | 87.5 | 95.5 | 91.2 | 85.9 | 83.9 | 71.8 | 75.4 |
SVLC-R | 58.8 | 79.4 | 91.2 | 91.3 | 81.2 | 64.2 | 68.6 | 69.1 |
SVLC-R+L | 59.7 | 75.8 | 81.1 | 88.1 | 76.8 | 62.6 | 64.1 | 66.7 |
CoN-CLIP | 63.2 | 87.9 | 78.0 | 91.8 | 81.0 | 66.3 | 63.7 | 67.0 |
TripletCLIP | 54.7 | 87.3 | 85.8 | 94.4 | 86.7 | 82.8 | 66.5 | 72.7 |
CLIC-COCO | 62.7 | 88.8 | 82.4 | 93.6 | 85.9 | 71.6 | 65.7 | 72.8 |
CLIC-LAION | 61.7 | 81.6 | 87.4 | 93.4 | 86.3 | 72.4 | 73.1 | 74.2 |
CLIC-RedCaps | 62.3 | 84.2 | 88.7 | 93.3 | 86.9 | 74.0 | 71.8 | 73.3 |
Downstream Evaluations | Sugarcrepe++ | Sugarcrepe | ||||||||||
Method | Classification | COCO Ret. | Rep. | Swap | Add | Rep. | Swap | |||||
IMNET | ZS-10 | Text | Image | ITT | TOT | ITT | TOT | ITT | ITT | ITT | ||
CLIP [40] | 63.3 | 61.4 | 74.1 | 54.6 | 69.5 | 60.5 | 45.7 | 25.9 | 72.9 | 80.0 | 62.7 | |
NegCLIP† | 61.0 | 60.5 | 72.0 | 59.7 | 67.9 | 64.7 | 54.5 | 36.1 | 82.4 | 80.9 | 70.7 | |
Single image Baselines (Non-Concat) | ||||||||||||
Fine-tuned | 62.1 | 61.8 | 76.4 | 60.4 | 69.6 | 69.2 | 49.8 | 35.9 | 83.4 | 80.7 | 68.8 | |
+ Single hard-negative | 60.6 | 59.8 | 70.4 | 57.9 | 66.6 | 63.8 | 55.0 | 36.2 | 81.9 | 80.5 | 70.6 | |
+ Multi Positives | 58.5 | 59.2 | 58.2 | 56.6 | 70.2 | 48.5 | 58.5 | 23.1 | 77.7 | 76.0 | 64.7 | |
Single-Img Baseline | 61.3 | 60.6 | 67.3 | 58.0 | 74.1 | 53.9 | 60.2 | 27.8 | 81.0 | 79.9 | 69.7 | |
Ablating CLIC (Concat) | ||||||||||||
Fine-tuned Concat | 62.6 | 61.2 | 76.5 | 59.1 | 69.8 | 66.7 | 49.3 | 35.9 | 82.7 | 81.9 | 68.2 | |
+Single hard-negative | 60.9 | 60.0 | 73.9 | 57.8 | 68.7 | 63.6 | 53.7 | 34.2 | 84.2 | 83.8 | 73.2 | |
+Multi Positives | 61.3 | 59.6 | 61.3 | 54.8 | 74.2 | 47.4 | 59.0 | 19.5 | 84.2 | 81.9 | 70.9 | |
+Uni-Modal Loss | 60.9 | 59.3 | 59.9 | 54.5 | 74.2 | 50.7 | 59.1 | 20.3 | 83.4 | 81.6 | 70.3 | |
+CLIP iterate (CLIC) | 61.7 | 61.0 | 75.9 | 60.0 | 75.6 | 60.1 | 61.1 | 27.9 | 84.5 | 84.0 | 73.7 |
Downstream Evaluations | SugarCrepe++ | SugarCrepe | ||||||||||||
Fine-tuning | Classification | Text Ret. | Image Ret. | Replace | Swap | Add | Rep. | Swap | ||||||
Method | Epochs | IMNET | ZS-10 | COCO | F30k | COCO | F30k | ITT | TOT | ITT | TOT | ITT | ITT | ITT |
CLIP [40] | - | 63.3 | 61.4 | 74.1 | 95.1 | 54.6 | 83.5 | 69.5 | 60.5 | 45.7 | 25.9 | 72.9 | 80.0 | 62.7 |
CLIC-COCO∗ | 2 | 63.2 | 60.9 | 81.7 | 96.6 | 66.3 | 89.8 | 73.0 | 78.3 | 53.0 | 45.4 | 85.6 | 83.3 | 69.9 |
CLIC-COCO∗ | 3 | 62.9 | 60.9 | 82.3 | 97.1 | 67.5 | 90.3 | 73.5 | 78.7 | 52.6 | 45.8 | 85.2 | 83.5 | 70.1 |
CLIC-COCO∗ | 5 | 62.7 | 60.7 | 82.9 | 97.1 | 68.2 | 90.2 | 73.5 | 79.4 | 52.9 | 47.2 | 85.6 | 83.7 | 69.3 |
Downstream Evaluations | SugarCrepe++ | SugarCrepe | ||||||||||||
Method | Frozen | Classification | Text Ret. | Image Ret. | Repl. | Swap | Add | Repl. | Swap | |||||
IMNET | ZS-10 | COCO | F30k | COCO | F30k | ITT | TOT | ITT | TOT | ITT | ITT | ITT | ||
CLIP [40] | 63.3 | 61.4 | 74.1 | 95.1 | 54.6 | 83.5 | 69.5 | 60.5 | 45.7 | 25.9 | 72.9 | 80.0 | 62.7 | |
CLIC | None | 60.5 | 60.8 | 77.7 | 94.7 | 62.1 | 87.4 | 72.2 | 60.0 | 60.0 | 25.1 | 84.5 | 83.6 | 71.9 |
CLIC | Text | 60.2 | 60.3 | 75.8 | 93.4 | 61.1 | 86.5 | 72.5 | 60.5 | 49.8 | 25.9 | 76.8 | 80.1 | 64.4 |
CLIC | Vision | 61.7 | 61.0 | 75.9 | 95.0 | 60.0 | 86.7 | 75.6 | 60.1 | 61.1 | 27.9 | 84.5 | 84.0 | 73.7 |
Downstream Evaluations | SugarCrepe++ | SugarCrepe | ||||||||||||
Method | Positives | Classification | Text Ret. | Image Ret. | Replace | Swap | Add | Replace | Swap | |||||
IMNET | ZS-10 | COCO | F30k | COCO | F30k | ITT | TOT | ITT | TOT | ITT | ITT | ITT | ||
CLIP [40] | 63.3 | 61.4 | 74.1 | 95.1 | 54.6 | 83.5 | 69.5 | 60.5 | 45.7 | 25.9 | 72.9 | 80.0 | 62.7 | |
CLIC | 62.0 | 61.3 | 75.5 | 94.4 | 60.0 | 86.4 | 69.0 | 67.1 | 51.8 | 35.8 | 85.9 | 83.1 | 70.9 | |
CLIC | 61.8 | 61.2 | 75.5 | 94.6 | 59.9 | 86.5 | 69.7 | 68.6 | 53.3 | 35.3 | 85.6 | 82.8 | 72.3 | |
CLIC | 61.7 | 60.8 | 75.8 | 94.9 | 59.8 | 86.6 | 73.8 | 61.7 | 58.7 | 30.2 | 84.5 | 83.8 | 72.6 | |
CLIC | 61.7 | 61.0 | 75.9 | 95.0 | 60.0 | 86.7 | 75.6 | 60.1 | 61.1 | 27.9 | 84.5 | 84.0 | 73.7 | |
CLIC | 61.7 | 60.7 | 76.2 | 95.1 | 59.8 | 86.7 | 75.6 | 59.3 | 61.5 | 27.0 | 84.1 | 83.6 | 73.0 |
Model | Source | Key |
ViT-B/32 | OpenClip | openai |
ViT-B/16 | OpenClip | openai |
ViT-L/14 | OpenClip | openai |
NegCLIP | GitHub | mertyg/vision-language-models-are-bows |
DAC-SAM | GitHub | SivanDoveh/DAC |
SVLC-R+L | GitHub | SivanDoveh/TSVLC |
CoN-CLIP | GitHub | jaisidhsingh/CoN-CLIP |
TripletCLIP | GitHub | tripletclip/TripletCLIP |
EVA-CLIP | OpenClip | eva02_large_patch14_clip_224 |
CLIPA | HF-Hub | UCSC-VLAA / hf-hub:UCSC-VLAA/ViT-L-14-CLIPA-datacomp1B |
CLIPS | HF-Hub | UCSC-VLAA / ViT-L-14-CLIPS-224-Recap-DataComp-1B |
Model | Source | Food101 | Cifar10 | Cifar100 | Cars | FGVS | DTD | Pets | Caltech101 | Flowers | Country211 | Mean |
CLIP | [40] | 93.6 | 96.4 | 76.3 | 76.5 | 30.4 | 54.6 | 94.9 | 86.4 | 80.6 | 33.1 | 72.3 |
CoN-CLIP | [43] | 93.9 | 96.0 | 79.4 | 76.3 | 30.4 | 56.8 | 93.6 | 86.8 | 78.1 | 33.0 | 72.4 |
CLIC-LAION | ours | 93.3 | 96.4 | 78.3 | 75.0 | 34.1 | 56.5 | 89.6 | 86.6 | 75.1 | 33.6 | 71.8 |
CLIC-RedCaps | ours | 92.5 | 96.3 | 78.0 | 76.0 | 30.8 | 56.6 | 91.7 | 87.3 | 78.6 | 32.9 | 72.1 |
CLIPA | [24] | 94.9 | 98.7 | 88.7 | 92.7 | 42.2 | 69.5 | 95.0 | 88.5 | 81.8 | 31.3 | 78.3 |
CLIC-LAION | ours | 93.9 | 98.8 | 88.4 | 92.9 | 43.1 | 67.8 | 95.7 | 88.4 | 80.7 | 30.8 | 78.0 |
CLIC-RedCaps | ours | 94.5 | 98.4 | 88.0 | 93.3 | 41.9 | 69.5 | 95.9 | 89.3 | 80.6 | 31.5 | 78.3 |
EVA-CLIP | [45] | 94.1 | 99.7 | 90.6 | 89.7 | 37.4 | 61.8 | 94.7 | 88.6 | 77.0 | 32.7 | 76.6 |
CLIC-LAION | ours | 94.1 | 99.7 | 90.6 | 89.7 | 37.4 | 61.8 | 94.7 | 88.6 | 77.0 | 32.7 | 76.6 |
CLIC-RedCaps | ours | 92.9 | 99.4 | 89.9 | 88.8 | 33.4 | 62.6 | 92.3 | 89.1 | 77.4 | 32.5 | 75.8 |
CLIPS | [30] | 93.2 | 98.2 | 86.9 | 91.6 | 39.4 | 66.2 | 94.9 | 86.9 | 78.2 | 29.4 | 76.5 |
CLIC-LAION | ours | 91.8 | 98.1 | 87.1 | 90.6 | 37.6 | 64.2 | 93.8 | 86.8 | 78.1 | 27.2 | 75.5 |
CLIC-RedCaps | ours | 92.7 | 98.0 | 86.8 | 91.2 | 37.7 | 67.1 | 94.5 | 87.8 | 80.8 | 30.0 | 76.7 |
ZS-Class. | Text-Ret. | Image-Ret. | |||||
Model | Source | IMNET | Avg. | COCO | F30k | COCO | F30k |
CLIP | [40] | 68.3 | 65.5 | 76.2 | 96.4 | 57.6 | 85.6 |
NegCLIP† | Ours | 64.5 | 64.0 | 74.2 | 94.7 | 65.8 | 89.7 |
CoN-CLIP∗ | [43] | 68.9 | 69.1 | 73.9 | 93.5 | 60.1 | 88.1 |
CLIC-LAION | Ours | 66.5 | 62.3 | 77.6 | 96.9 | 63.6 | 88.9 |
CLIC-RedCaps | Ours | 67.3 | 64.4 | 79.0 | 97.0 | 62.0 | 88.0 |
Compositionality: SugarCrepe++ | ||||||||
Model | Source | Rep-obj | Rep-Att | Rep-Rel | Swap-obj | Swap-Att | Avg-ITT | Avg-TOT |
CLIP | [40] | 89.6 | 67.5 | 53.1 | 39.6 | 48.3 | 59.6 | 45.7 |
NegCLIP† | Ours | 88.7 | 66.0 | 46.9 | 47.3 | 53.7 | 60.5 | 51.9 |
CoN-CLIP | [43] | 88.1 | 66.6 | 52.3 | 44.1 | 51.9 | 60.5 | 55.5 |
CLIC-LAION | Ours | 91.6 | 75.9 | 62.5 | 55.9 | 62.0 | 69.6 | 45.7 |
CLIC-RedCaps | Ours | 91.5 | 76.0 | 64.5 | 55.9 | 61.1 | 69.8 | 38.6 |
Compositionality: SugarCrepe | ||||||||
Add-obj | Add-Att | Rep-Obj | Rep-Att | Rep-Rel | Swap-Obj | Swap-Att | ||
CLIP | [40] | 78.4 | 66.8 | 93.5 | 81.0 | 66.6 | 60.0 | 65.0 |
NegCLIP† | Ours | 80.0 | 87.3 | 93.9 | 82.2 | 68.1 | 69.0 | 70.1 |
CoN-CLIP | [43] | 87.3 | 79.6 | 93.6 | 81.0 | 53.3 | 59.2 | 65.2 |
CLIC-LAION | Ours | 88.4 | 91.0 | 95.6 | 86.5 | 75.5 | 71.0 | 73.9 |
CLIC-RedCaps | Ours | 87.2 | 88.6 | 94.9 | 85.7 | 73.3 | 69.0 | 71.3 |
Compositionality: SugarCrepe++ | ||||||||
Model | Source | Rep-obj | Rep-Att | Rep-Rel | Swap-obj | Swap-Att | Avg-ITT | Avg-TOT |
CLIP | [40] | 90.6 | 67.5 | 54.0 | 43.6 | 45.6 | 60.3 | 50.3 |
NegCLIP† | Ours | 90.5 | 68.9 | 52.3 | 46.5 | 54.5 | 62.5 | 52.4 |
CoN-CLIP | [43] | 92.2 | 68.6 | 55.6 | 44.1 | 47.9 | 61.7 | 54.9 |
CLIC-LAION | Ours | 94.6 | 79.6 | 62.8 | 59.6 | 60.0 | 71.3 | 46.1 |
CLIC-RedCaps | Ours | 93.5 | 75.0 | 59.5 | 54.7 | 55.7 | 67.7 | 41.3 |
Compositionality: SugarCrepe | ||||||||
Add-obj | Add-Att | Rep-Obj | Rep-Att | Rep-Rel | Swap-Obj | Swap-Att | ||
CLIP | [40] | 78.3 | 71.5 | 94.1 | 79.2 | 65.1 | 60.4 | 62.3 |
NegCLIP† | Ours | 81.2 | 86.0 | 94.3 | 83.1 | 71.1 | 69.4 | 70.6 |
CoN-CLIP | [43] | 90.2 | 77.6 | 95.2 | 81.7 | 67.0 | 65.3 | 63.1 |
CLIC-LAION | Ours | 86.1 | 89.3 | 96.5 | 86.4 | 73.8 | 71.0 | 71.6 |
CLIC-RedCaps | Ours | 85.8 | 82.8 | 95.4 | 82.9 | 70.0 | 64.5 | 67.7 |
Compositionality: SugarCrepe++ | ||||||||
Model | Source | Rep-obj | Rep-Att | Rep-Rel | Swap-obj | Swap-Att | Avg-ITT | Avg-TOT |
CLIPA | [24] | 94.5 | 73.6 | 56.3 | 41.2 | 50.4 | 63.2 | 57.2 |
CLIC-LAION | Ours | 93.9 | 79.1 | 62.1 | 52.2 | 57.9 | 69.1 | 52.1 |
CLIC-RedCaps | Ours | 95.0 | 80.6 | 64.8 | 60.4 | 62.2 | 72.6 | 46.8 |
[…]
CLIC-LAION | Ours | 93.8 | 79.7 | 61.8 | 55.9 | 61.1 | 70.4 | 45.2 |
CLIC-RedCaps | Ours | 95.2 | 80.7 | 63.8 | 59.2 | 60.0 | 71.8 | 41.9 |
[…]
CLIC-LAION | Ours | 95.5 | 84.8 | 72.6 | 66.1 | 80.0 | 79.8 | 62.3 |
CLIC-RedCaps | Ours | 95.9 | 85.1 | 73.7 | 68.6 | 81.7 | 81.0 | 62.9 |
Compositionality: SugarCrepe | ||||||||
Add-obj | Add-Att | Rep-Obj | Rep-Att | Rep-Rel | Swap-Obj | Swap-Att | ||
CLIPA | [24] | 88.8 | 82.6 | 96.8 | 83.0 | 68.8 | 62.4 | 65.1 |
CLIC-LAION | Ours | 93.5 | 92.2 | 97.1 | 89.3 | 78.3 | 72.2 | 70.9 |
CLIC-RedCaps | Ours | 94.5 | 92.3 | 97.0 | 86.9 | 73.8 | 69.8 | 73.3 |
[…]
CLIC-LAION | Ours | 85.5 | 91.3 | 96.4 | 86.5 | 72.7 | 70.6 | 70.5 |
CLIC-RedCaps | Ours | 90.7 | 90.6 | 96.7 | 86.7 | 73.2 | 70.2 | 70.3 |
[…]
CLIC-LAION | Ours | 96.6 | 96.8 | 98.7 | 93.0 | 85.5 | 82.0 | 89.3 |
CLIC-RedCaps | Ours | 95.1 | 95.5 | 97.9 | 92.2 | 83.2 | 81.2 | 87.8 |
Retrieval | WinoGround | ||||||
Method | Text | Image | |||||
COCO | F30k | COCO | F30k | Text | Image | Group | |
CLIP [40] | 56.0 | 85.1 | 35.2 | 64.7 | 28.7 | 11.0 | 8.5 |
CoN-CLIP∗ [43] | 55.8 | 82.5 | 37.8 | 69.3 | 29.5 | 11.0 | 8.0 |
CLIC-LAION | 56.8 | 85.4 | 40.2 | 70.4 | 30.2 | 10.7 | 8.2 |
CLIC-RedCaps | 58.3 | 86.9 | 39.0 | 68.8 | 32.2 | 12.2 | 9.0 |
CLIPA | 64.4 | 90.4 | 46.2 | 73.8 | 32.7 | 8.0 | 7.0 |
CLIC-LAION | 63.0 | 86.7 | 46.8 | 73.6 | 28.7 | 8.7 | 6.7 |
CLIC-RedCaps | 63.2 | 89.1 | 46.6 | 73.9 | 32.0 | 11.2 | 7.2 |
EVA-CLIP | 63.7 | 90.5 | 46.8 | 77.3 | 32.7 | 12.5 | 10.2 |
CLIC-LAION | 63.6 | 90.5 | 46.5 | 77.4 | 31.7 | 12.7 | 8.7 |
CLIC-RedCaps | 63.2 | 89.8 | 46.4 | 76.7 | 36.0 | 11.7 | 8.2 |
CLIPS | 73.6 | 95.7 | 54.2 | 82.6 | 36.2 | 16.0 | 12.5 |
CLIC-LAION | 71.0 | 95.3 | 56.6 | 85.7 | 37.5 | 16.0 | 13.2 |
CLIC-RedCaps | 74.6 | 96.3 | 56.6 | 84.3 | 41.7 | 17.5 | 15.2 |
Retrieval | ||||
Method | Text | Image | ||
COCO | F30k | COCO | F30k | |
CLIP [40] | 49.5 | 78.5 | 30.1 | 59.1 |
NegCLIP∗ [49] | 60.8 | 85.1 | 45.1 | 70.9 |
DAC-LLM [8] | 33.0 | 64.9 | 36.7 | 53.2 |
DAC-SAM [8] | 33.0 | 61.5 | 33.2 | 58.4 |
SVLC-R [9] | 45.0 | 73.5 | 35.4 | 64.5 |
SVLC-R+L [9] | 42.8 | 70.4 | 35.6 | 63.7 |
CoN-CLIP∗ [43] | 48.5 | 75.7 | 30.3 | 59.7 |
TripletCLIP [37] | 46.9 | 73.7 | 32.3 | 60.2 |
CLIC-COCO∗ | 60.0 | 84.7 | 41.1 | 68.3 |
CLIC-LAION | 52.2 | 79.9 | 35.2 | 62.6 |
CLIC-RedCaps | 52.5 | 79.4 | 33.8 | 62.8 |